OpenStudy (anonymous):

a student open a mathematics book to two facing pages. The product of the page numbers is 702. Find the page numbers.

OpenStudy (goformit100):

@satellite73

OpenStudy (saifoo.khan):

26 and 27

OpenStudy (anonymous):

Since the numbers must be consecutive, we can call them \(n\) and \(n+1\). Thus, we know\(n(n+1)=702\). Now, you can distribute and solve for \(n\) and thus get \(n+1\) as well.

OpenStudy (saifoo.khan):

\[\sqrt{702} = 26.49\] so, 26 and 27. my shortcut. :D

OpenStudy (anonymous):

oh heck guess and check for this one bet you get it on the first or second try

OpenStudy (anonymous):

That's a good method too, though not quite as kosher. :P Sorry, that equation above was a little hard to read in my post. Here it is again: \(n(n+1)=702\)

OpenStudy (anonymous):

\(25\times 26=650\) nope

OpenStudy (anonymous):

\(26\times 27=702\) ok that works

OpenStudy (saifoo.khan):

But we have to take the value after the integer. Like 26.49 so 0.49 is bigger than 26. therefore 26 and the integer after it.

OpenStudy (saifoo.khan):

@satellite73 , Even single shot will give us the answer. ;D

OpenStudy (anonymous):

@saifoo.khan i had a single shot earlier then a double

OpenStudy (saifoo.khan):

@satellite73 Lol. but now i told you the trick for success in first shot. :D